### v2

# Undersmoothed LASSO Models for Propensity Score Weighting and Synthetic Negative Control Exposures for Bias Detection

Abstract
The propensity score (PS) is often used to control for large numbers of covariates in high-dimensional healthcare database studies. The least absolute shrinkage and selection operator (LASSO) is a data-adaptive prediction algorithm that has become the most widely used tool for large-scale PS estimation in these settings. However, recent work has shown that the use of data-adaptive algorithms for PS estimation can come at the cost of slow convergence rates, resulting in PS-based causal estimators having poor statistical properties. While this can create challenges for the use of data-driven algorithms for PS analyses, both theory and simulations have shown that LASSO PS models can converge at a fast enough rate to provide asymptotically efficient PS weighted causal estimators. In order to achieve asymptotic efficiency, however, LASSO PS weighted estimators need to be properly tuned, which requires undersmoothing the fitted LASSO model. In this paper, we discuss challenges in determining how to undersmooth LASSO models for PS weighting and consider the use of balance diagnostics to select the degree of undersmoothing. Because no tuning criteria is universally best, we propose using synthetically generated negative control exposure studies to detect bias across alternative analytic choices. Specifically, we show that synthetic negative control exposures can identify undersmoothing techniques that likely violate partial exchangeability due to lack of control for measured confounding. We use a series of numerical studies to investigate the performance of alternative balance criteria to undersmooth LASSO PS-weighted estimators, and the use of synthetic negative control exposure studies to detect biased analyses.
1 Introduction
Routinely collected healthcare data, including administrative claims and electronic health records, are increasingly being used to generate real‐world evidence on the effects of medical products. However, confounding stemming from nonrandomized exposures remains a fundamental obstacle to effectively utilizing these data sources for real-world evidence generation. To improve confounding control in healthcare database studies, data‐driven algorithms can be used to leverage the large volume of information in these data sources to identify features that indirectly capture information on unspecified confounding factors [1, 2, 3, 4]. A large literature has shown that supplementing investigator-specified covariates with large numbers of empirically identified features can often improve confounding control compared with adjustment for investigator-specified covariates alone [1, 2, 3, 4, 5]. In healthcare database studies, however, outcome events are often rare, making it difficult to empirically identify and model outcome associations for large numbers of covariates. Consequently, the propensity score (PS), defined as the conditional probability of exposure given a set of covariates, is widely used for large-scale covariate adjustment in these settings [6, 7].
Fitting large-scale PS models can capture more confounding factors, but it can also result in reduced covariate overlap, which can harm the properties of effect estimates and even violate positivity assumptions [8, 9]. Consequently, when fitting large-scale PS models, covariate selection is necessary to avoid problems of nonoverlap. The least absolute shrinkage and selection operator (LASSO) is a data-adaptive prediction algorithm that has become the most widely used tool for fitting large-scale PS models [4, 10]. LASSO uses penalized regression to data-adaptively shrink imprecise coefficients toward zero, effectively performing covariate selection to exclude less important predictors. However, recent work has shown that the use of data-adaptive machine-learning algorithms for estimating the PS can come at the cost of slow convergence rates. Specifically, it has been shown that if the PS model does not converge at a minimum rate of $n^{-1/4}$ (where $n$ is the sample size), then this can result in PS-based causal estimators having poor statistical properties with performance deteriorating as the dimension of the data increases (the curse of dimensionality) [11, 12, 13].
In recent work, Benkeser & van der Laan [14] showed that if the true regression function can be expressed as a linear combination of some expanded basis of the covariates (discussed further in Section 2), then a LASSO regression that is fitted on the expanded set of basis functions can estimate the true model at a rate of $n^{-1/3}$. Benkesser & van der Laan [14] further explain that since a very general class of functions can be expressed as a linear combination of an expanded set of indicator basis functions, LASSO regression is generally applicable for estimating nuisance functions at a fast enough rate for causal estimators to have desirable properties. Building on this work, Ertefaie et al [15] showed that LASSO models can provide asymptotically linear PS weighted estimators with variance converging to the nonparametric efficiency bound.
To achieve asymptotic efficiency, however, LASSO PS weighted estimators need to be properly tuned, which requires undersmoothing the fitted LASSO model (discussed further in Section 2). Ertefaie et al [15] showed that the degree of undersmoothing can be derived from the form of the target causal parameter’s efficient influence function, if that function is available. In practice, however, the efficient influence function is often not known or difficult to derive. Even when the analytic form of the efficient influence function is known, using the efficient influence function to undersmooth the LASSO model requires modeling the conditional outcome mean [15], which can be difficult in studies with rare outcome events. It remains unclear how to best tune LASSO PS weighted estimators when the efficient influence function is difficult to derive or unknown, or when modeling the conditional outcome mean is challenging.
In this work, we discuss challenges in determining how to undersmooth LASSO models for PS weighting and consider the use of balance diagnostics to select the degree of undersmoothing. Because no tuning criteria is universally best, we propose a framework to generate synthetic negative control exposure studies for bias detection. We show that if a given LASSO PS weighted analysis does not result in conditional independence between the synthetic exposures and observed outcome within unexposed individuals, then the same analysis is unlikely to satisfy partial exchangeability when applied to the original study population. Finally, we use a series of numerical studies to evaluate the performance of alternative undersmoothing criteria and the use of synthetic negative control exposure studies to detect biased analyses.
2 Methods
2.1 Framework & Setup
We assume that we have a sample of $n$ independent and identically distributed observations, $O_{1},O_{2},\cdots,O_{n}$, with data structure $O=\{Y,A,X\}$ drawn from a probability distribution $P(Y,A,X)$. In this data structure, $X$ is a $d$-dimensional vector of baseline covariates, $A$ is a binary exposure, and $Y$ is the observed outcome. Following Neyman [16] and Rubin [17], we define the effect of $A$ on $Y$ in terms of potential outcomes, $Y^{a=1}$ and $Y^{a=0}$, where the observed outcome, $Y$, corresponds with either $Y^{a=1}$ or $Y^{a=0}$ depending on whether the individual was exposed ($a=1$) or not exposed ($a=0$). Furthermore, let $P(Y^{a=0},A,X)$ represent the probability distribution for the counterfactual population under no exposure, and let $e(X)=P(A|X)$ represent the conditional probability of exposure given $X$ (i.e., the PS).
We assume that exposure is conditionally exchangeable given $X$, written as
$(Y^{a=1},Y^{a=0})\perp\!\!\!\perp A|X,$
[…]
$Y^{a=0}\perp\!\!\!\perp A|X.$
Partial exchangeability is a weaker condition than full conditional exchangeability. If an adjustment set fails to satisfy partial exchangeability, then the same adjustment set would also fail to satisfy full conditional exchangeability [18].
We further assume positivity and consistency [19]. Positivity, formally written as $10<e(X)<$ 1, implies that the true PS function, $e(X)$, is bounded away from 0 and 1. Consistency implies that the observed outcome for each individual is equal to the potential outcome under their observed exposure status, written as $Y^{a}=Y$ for $A=a$.
In addition to the standard assumptions for causal inference described above, we assume that the logit of the propensity score function, $logit(e(X))$, is linear in $X$, written as:
$logit(e(X))=\beta_{0}+X^{\top}\beta$ (1)
where $\beta$ is a $d$-dimensional vector of parameters and $\beta_{0}$ is a scalar. Healthcare database studies often consist of adjusting for a high-dimensional set sparse binary covariates where linearity assumptions in the PS model are often reasonable [1, 2, 3, 4, 5]. When the assumption of linearity in the relationship between $X$ and the logit of the propensity score is not reasonable, Benkeser & van der Laan [14] showed that under mild global smoothness assumptions, the baseline covariates, $X$, can be expanded into a series of $n(2^{d}-1)$ binary indicator variables (i.e., indicator basis functions), $W$, such that as $n\rightarrow\infty$ the logit of the propensity score function, $logit(e(X))$, can be approximated arbitrarily well by a linear combination of the indicator basis functions written as:
$logit(g(W))=\gamma_{0}+W^{\top}\gamma$ (2)
where $g(W)=P(A|W)$, $\gamma$ is a $n(2^{d}-1)$ dimensional vector of parameters, and $\gamma_{0}$ is a scalar. For a theoretical explanation on the construction of $W$, see Benkeser and van der Laan [14].
Throughout this paper, we focus on the application of LASSO models to the baseline covariates, $X$, under the assumption of linearity. However, the methods considered here are generally applicable to nonlinear settings by fitting a LASSO regression on a linear combination of the expanded indicator basis functions, $W$. This has been termed the Highly Adaptive Lasso—a nonparametric machine-learning prediction algorithm that has been shown to have theoretical guarantees on fast convergence rates under mild assumptions [14, 15]. While our focus is not on the Highly Adaptive Lasso, we do consider the application of the Highly Adaptive Lasso in some numerical studies in Section 3. For more on the Highly Adaptive Lasso, see Benkeser & van der Laan [14]. A general overview of the Highly Adaptive Lasso is provided in the Supplemental Appendix.
2.2 Undersmoothing LASSO PS Models
When using LASSO regression to estimate $e(X)$, the parameter vector, $\beta$, in Equation (1) is estimated based on the following penalized likelihood [4]:
${\cal L}(\beta)=\sum_{i=1}^{n}-A_{i}log(p(A_{i}|X_{i};\beta))-(1-A_{i})log(1-p(A_{i}|X_{i};\beta))+\lambda\sum_{j=1}^{d}\mid\beta_{j}\mid$ (3)
where $\lambda$ is the regularization tuning parameter, $n$ is the sample size, $d$ is the number of parameters in $\beta$, and $p(A_{i}|X_{i};\beta)$ follows the logistic model defined in Equation 1.
If the regularization tuning parameter, $\lambda$, in Equation 3 is chosen appropriately, then as $n\rightarrow\infty$, $\lambda$ will tend to zero and the coefficients within the LASSO model will converge to those in Equation 1. In finite samples, however, LASSO is just an approximation to $e(X)$ and different choices for $\lambda$ provide different LASSO estimators. The optimal choice for $\lambda$ depends on the purpose for which the LASSO model is used. If the goal is to optimize out-of-sample prediction, then $\lambda$ is often chosen using cross-validation, which we will refer to as $\lambda_{CV}$. Ertefaie et al [15] showed, however, that when using LASSO models for PS weighting, less regularization is needed to minimize bias in PS weighted estimators. This is referred to as undersmoothing the LASSO model and corresponds to choosing a value for $\lambda$ that is less than $\lambda_{CV}$ so that the model selects more covariates from $X$ compared to the model using $\lambda_{CV}$.
Undersmoothing can capture more confounder information by including more covariates from $X$ in the model. However, undersmoothing can also harm the overall accuracy of the fitted model by modeling spurious associations in the data (overfitting). Previous work has shown that cross-fitting or using the out-of-fold predictions from the LASSO model can improve properties of effect estimates by reducing problems caused by overfitting [12, 20]. Still, there is a tradeoff with undersmoothing LASSO models for PS estimation. Proper undersmoothing improves efficiency of causal estimators by controlling for more confounder information, but too much undersmoothing will eventually harm the accuracy of the estimated coefficients to a degree where the benefit of including more information in $X$ is outweighed by the cost of a poorly fit model producing unstable predictions [21].
Criteria have been proposed to help investigators properly undersmooth LASSO PS models when using estimators that target populations that are well defined with efficient influence functions that are easy to derive (e.g., inverse probability weighting to estimate the average treatment effect) [15, 22]. However, it remains unclear how to properly undersmooth LASSO PS models when the efficient influence function for the target parameter is difficult to derive or when the target population is not well defined. Here, we consider using balance metrics as a simple and general approach to undersmooth LASSO models for PS weighting.
2.3 Using Balance Metrics to Undersmooth LASSO PS Models
Metrics for evaluating covariate balance across exposure groups have become standard when evaluating PS models for confounding control [23, 24, 25]. Balance diagnostics have the benefit of being easy to implement and generally applicable. Still, it can be challenging to measure balance on the joint covariate distribution in high-dimensional settings [26]. If $X$ is binary and linearly related to exposure, however, assessing balance on the joint correlation structure is straightforward since balance on the marginal prevalence of the covariates implies balance on the full joint covariate distribution. As discussed previously, large-scale PS analyses in healthcare database studies usually consist of adjusting for a high-dimensional set of sparse binary covariates where linearity assumptions are often reasonable.
Here, we consider two commonly used balance metrics for such settings [24, 25]. Each metric uses the standardized difference to assess balance after PS weighting. Letting $\widehat{p}_{k,exposed}$ and $\widehat{p}_{k,unexposed}$ represent the sample prevalence for each binary covariate, $k$, in the exposed and unexposed groups, respectively, the standardized difference and balance metrics are defined as:
$s_{k}=\frac{(\widehat{p}_{k,exposed}-\widehat{p}_{k,unexposed})}{\sqrt{\frac{\widehat{p}_{k,exposed}(1-\widehat{p}_{k,exposed})+\widehat{p}_{k,unexposed}(1-\widehat{p}_{k,unexposed})}{2}}}$
  • Largest standardized absolute mean difference. Selects the lambda tuning parameter that minimizes the following after PS weighting: $max[|s_{1}|,|s_{2}|,...,|s_{d}|]$.
  • Average standardized absolute mean difference. Selects the lambda tuning parameter that minimizes the following after PS weighting: $\frac{1}{d}\sum_{k=1}^{d}|s_{k}|$.
When $X$ is not binary, or it is not reasonable to assume linearity in the association between $X$ and the log-odds of exposure, the above metrics are still generally applicable by replacing $X$ with an expanded set of indicator basis functions, $W$, as discussed previously. Since the log-odds of exposure is linear in $W$, it is conjectured that balance on the marginal prevalence of the generated binary indicators implies balance on the joint distribution of the original covariates. It is important to note that for continuous covariates, it is common practice to only balance covariate means, which is restrictive. Balancing the expanded set of indicator basis functions for non-binary covariates allows for stronger, more nonparametric balance guarantees.
While many other balance metrics could be considered, our goal is not to perform an exhaustive evaluation of approaches that could be used to undersmooth LASSO PS weighted estimators. Here, we consider two commonly used balance metrics for evaluating large-scale PS models but emphasize that no single metric is universally best. The best metric will depend on properties of the given study as well as the chosen weighting approach. Consequently, it can be difficult to know which metric best correlates with bias for the study at hand and if the level of balance achieved is adequate to remove bias caused by measured confounders [26]. Therefore, to address this challenge, we suggest supplementing balance diagnostics by using synthetically generated negative control exposures to help detect biased analyses.
2.4 Using Synthetic Data for Bias Detection with LASSO PS Analyses
The use of real data to generate synthetic cohorts, where exposure-outcome associations are known by design and simulated patterns of confounding mimic the observed data structure, have become increasingly used to provide a benchmark for validating analytic choices for causal inference [27, 28, 29]. Simulation frameworks that use real data to generate synthetic cohorts have generally been termed ‘plasmode simulation’ [30, 31, 32, 33].
The use of plasmode simulation for model validation in causal inference has been compared to the use of cross-validation for prediction models [34]. Schuler et al [34] explain, however, that validation frameworks based on plasmode simulation are more limited since they are not ‘model free’; they require partial simulation of the data structure. This creates two fundamental challenges when using plasmode simulation to evaluate causal inference methods: 1) Advani et al [35] showed that if the simulation framework does not closely approximate the true data generating distribution, then the use of synthetically generated data as a diagnostic tool in causal inference can be misleading; 2) even when the simulation framework closely approximates the true data generating process, Schuler et al [34] warn that the use of plasmode simulation for model validation could still be biased towards favoring causal inference methods that mimic the modeling choices made when generating the synthetic datasets (overfitting to the synthetic data).
To mitigate the challenges outlined above when validating LASSO PS weighted analyses, we propose using synthetically generated negative control exposures. Frameworks for generating synthetic negative control exposures address the first challenge by not attempting to simulate cohorts that approximate the full data distribution, $P(Y,A,X)$. Simulating the full confounding structure can be difficult in studies where modeling the conditional outcome mean is challenging relative to the propensity score, which is our focus here. Instead, frameworks for generating synthetic negative control exposures only require a model for the exposure to approximate a simpler confounding structure that is related to the data distribution for the counterfactual population under no exposure, $P(Y^{a=0},A,X)$ (Discussed further in Sections 2.4.1 and 2.4.2).
Because the synthetic cohorts do not approximate the full data distribution, they are simply used to detect analyses that are unlikely to satisfy partial exchangeability (discussed in Section 2.4.1). Using synthetic cohorts for bias detection rather than model selection helps to mitigate the challenge of making analytic decisions that overfit to the synthetic data. Below, we formally define negative control exposures and describe how they relate to conditional exchangeability in the full study population. We then define and outline a framework to generate synthetic negative control exposures to detect bias in LASSO PS weighted analyses.
2.4.1 Negative Control Exposures
In addition to the data structure, $\{Y,X,A\}$, defined previously, assume we observe a binary random variable, $Z$, generated with probability $g(X)=P(Z|X)$. Let $Y^{z}$ represent the potential outcome under the condition $Z=z$, and $Y^{(a,z)}$ the potential outcome that would be observed under the condition $A=a$ and $Z=z$. We define $Z$ to be a negative control exposure if the following conditions hold: 1) there is no causal relationship between the negative control exposure and the outcome; 2) there is no unmeasured common cause between the negative control exposure and the outcome; 3) there is no association between the negative control exposure and the exposure conditional on $X$, and 4) the conditional distribution of $A$ given $X$ is structurally equivalent to the conditional distribution of $Z$ given $X$ in the sense that $e(X)$ and $g(X)$ contain the same information with respect to causal relationships between $X$ and $A$ and between $X$ and $Z$. This can be stated more formally in the following definition.
Definition 1.
Assume we have the data structure, $\{Y,X,A,Z\}$, where $Y$, $X$, and $A$ are defined in Section 2.1. Let $Z$ be a binary random variable generated with probability $g(X)=P(Z|X)$, where $10<g(X)<$ 1. We define $Z$ to be a negative control exposure if for all $a$ and $z$: 1) $Y^{(a,z)}=Y^{a}$; 2) $Y^{a}\perp\!\!\!\perp Z|X$; 3) $Z\perp\!\!\!\perp A|X$; and 4) $g(X)=f(e(X))$, where $f(\cdot)$ is any 1-to-1 function of $e(X)$, where $e(X)=P(A|X)$.
It is important to note that there are many alternative definitions for a negative control exposure [36]. Which definition is most appropriate depends on how the negative control exposure is used. Definition 1 follows similar conditions to those of a disconnected negative control defined in Kummerfeld et al [37] and Shi et al [36]. The difference being that Condition 4 in Definition 1 (structural equivalence) is written only in terms of measured covariates $X$ and assumes no unmeasured confounding. In addition, the definition we use here is more restrictive since the assumption of structural equivalence (Condition 4) requires that there is a unique mapping between $e(X)$ and $g(X)$. This ensures that $e(X)$ and $g(X)$ are balancing scores for both $A$ and $Z$ [6], which implies that the set of all covariates and higher order terms (including interactions) affecting exposure and the negative control exposure are equivalent. This condition implies, but is even stronger, than Markov Equivalence where two conditional probability distributions, $e(X)$ and $g(X)$, share the same causal graph with respect to conditional independencies between $X$ and $A$ and between $X$ and $Z$ [38, 39]. Figure 1 illustrates a directed acyclic graph for a negative control exposure that satisfies Definition 1.
$A$$Y$$Z$$X$
Figure 1: Directed acyclic graph illustrating possible causal relationships between a set of covariates ($X$), a negative control exposure ($Z$), an exposure ($A$) and an outcome ($Y$).
Definition 1 implies that if conditioning on a set of covariates is sufficient to satisfy exchangeability in the study population, then conditioning on the same set of covariates will satisfy conditional independence between the observed outcome, $Y$, and the negative control exposure, $Z$, in the study population. This is formally stated in the following proposition. The proof to Proposition 1 follows immediately from the definition of a negative control exposure (details are provided in the Supplemental Appendix).
Proposition 1.
Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure. If $Y^{a}\perp\!\!\!\perp A|X$ it follows that $Y\perp\!\!\!\perp Z|X$.
Proposition 1 implies that negative control exposures can be used for bias detection. If adjustment for a set of covariates, $X$, does not result in unbiased null effect estimates when evaluating the association between a negative control exposure and the observed outcome, then under the assumption of no unmeasured confounding (for both the exposure and negative control exposure), then the analytic approch does not sufficiently control for all confounder information in $X$ (e.g., a LASSO PS weighted analysis with improper undersmoothing). It is important to note that negative controls are typically used for bias detection of unmeasured confounding which requires additional assumptions on the causal structure of unmeasured confounders between exposure and the negative control exposure. Here, we are simply interested in the use of negative control exposures to detect bias caused by the lack of control for measured confounders under the assumption of no unmeasured confounding.
Although negative control exposures can be used for bias detection, Definition 1 makes no statement on how covariates in, $X$, relate to the negative control exposure in terms of strength and direction of effects. Therefore, it is difficult to know how the magnitude of bias caused by $X$ on the effect of $Z$ on $Y$ relates to the magnitude of bias caused by $X$ on the effect of $A$ on $Y$ without additional assumptions on how $g(X)$ relates to $e(X)$. Still, even if $g(X)=e(X)$ so that the distribution of covariates across the exposure groups match those across the negative control exposure groups, the structure and magnitude of bias caused by $X$ on the relationship between $A$ and $Y$ can differ substantially from that between $Z$ and $Y$ if $A$ is causally related to the outcome (since $Z$ has a backdoor path to the outcome through $A$, where the only backdoor path from $A$ to $Y$ is through $X$). If one conditions on $A$, the magnitude of bias caused by $X$ on the relationship between $A$ and $Y$ versus that between $Z$ and $Y$ can still differ if covariate effects on the outcome vary between strata of $A$ [40].
If one wishes to avoid making assumptions on the causal relationship between $A$ and $Y$ when using negative control exposures to understand the structure and magnitude of confounding bias, one could restrict on $A=0$ so that the confounding structure between $Z$ and $Y$ among the unexposed approximates the confounding structure between $A$ and the counterfactual outcome, $Y^{a=0}$, in the study population. After restricting to the unexposed population a similar result to Proposition 1 holds. Specifically, it can be shown that conditional independence between a negative control exposure, $Z$, and the observed outcome, $Y$, within the unexposed population is connected to conditional independence between the actual exposure, $A$, and counterfactual outcome, $Y^{a=0}$, in the full study population (i.e., partial exchangeability). This can be stated more formally in the following proposition. The proof to Proposition 2 is provided in the Supplemental Appendix.
Proposition 2.
Assume we have the data structure, $\{Y,X,A,Z\}$ defined previously, where $Z$ is a negative control exposure. If $Y^{a=0}\perp\!\!\!\perp A|X$ it follows that $Y\perp\!\!\!\perp Z|X,A=0$.
Proposition 2 implies that if adjusting for a set of variables, $X$, is sufficient to satisfy partial exchangeability in the full study population, then proper adjustment for the same set of variables will satisfy conditional independence between the observed outcome, $Y$, and the negative control exposure, $Z$, in the unexposed population. This result is similar to Proposition 1, but since we are restricting to the unexposed population, the negative control exposure can only be used to detect violations of partial exchangeability (as opposed to full conditional exchangeability).
As mentioned previously, the purpose for restricting to unexposed individuals when estimating the effect of $Z$ on $Y$, is to approximate the confounding structure between $A$ and the counterfactual outcome, $Y^{a=0}$, in the study population. Similarity between these two confounding structures depends only on how $g(X)$ relates to $e(X)$. If $Y^{a=0}$ could be observed in the full study population and $g(X)=e(X)$ so that the distribution of covariates across negative control exposure groups match those across the exposure groups, then it is straightforward that the magnitude of bias caused by $X$ on the effect of $Z$ on $Y^{a=0}$ will be equal the magnitude of bias caused by $X$ on the effect of $A$ on $Y^{a=0}$ [40].
In practice, however, $Y^{a=0}$ can only be observed for those where $A=0$. After restricting on $A=0$, it is not possible for the distribution of covariates across the negative control exposure groups to match those across the exposure groups in the full population. Still, as long as the odds of exposure are proportional to the odds of the negative control exposure, general patterns in the separation of covariates across exposure groups will be similar to those across the negative control exposure groups after restricting to the unexposed (i.e., restricting on $A=0$). We define a negative control exposure with proportional odds as follows:
Definition 2.
Assume we have the data structure, $\{Y,X,A,Z\}$, where $Y$, $X$, and $A$ are defined in Section 2.1, and $Z$ is a binary random variable that satisfies the conditions for a negative control exposure as stated in Definition 1. We say that $Z$ is a negative control exposure with proportional odds if $\frac{g(X)}{1-g(X)}\propto\frac{e(X)}{1-e(X)}$.
Comment 1.
Definition 2 satisfies the conditions in Definition 1, but is more restrictive. Specifically, in addition to requiring a 1-to-1 mapping between $e(X)$ and $g(X)$, the proportional odds assumption places restrictions on how the strength and direction of covariate effects on exposure relate to the strength and direction of covariate effects on the negative control exposure. Specifically, if the $logit(e(X))$ follows Equation (1) (i.e. is linear in $X$), the proportional odds assumption ensures that the coefficient for each covariate (conditional effect) in the linear predictor for the $logit(e(X))$ and the $logit(g(X))$ is equivalent, with the only difference between the linear predictors potentially being the intercept.
Comment 2.
The goal of the proportional odds assumption is to have covariate differences across negative control exposure groups within the unexposed that approximate covariate differences across the exposure groups in the full population. If covariate differences across exposure groups match those across negative control exposure groups within the unexposed, then under the assumption that the association between $X$ and $A$ is linear, it can be shown that the magnitude of confounding bias caused by $X$ on the relationship between $A$ and $Y^{a=0}$, will be equal to the bias caused by $X$ on the relationship between $Z$ and the observed outcome, $Y$, among the unexposed on the absolute scale (see proof in Supplemental Appendix S1.4). If the relationship between $X$ and $A$ is not linear, but is linear in some expanded basis (e.g., the Highly Adaptive Lasso), then the same arguments hold and are generally applicable by replacing $X$ with the expanded set of basis functions.
Comment 3.
In practice, however, the proportional odds assumption only guarantees that covariate differences across negative control exposure groups within the unexposed will approximate differences across exposure groups in the full study population; they will not be exact. Still, even when covariate differences across negative control exposure groups within the unexposed do not mirror those of the full exposed and unexposed populations, we conjecture that the strength of confounding caused by covariates on the relationship between $Z$ and $Y$ among the unexposed will be approximately proportional to the strength of confounding caused by covariates on the relationship between $A$ and $Y^{a=0}$ in the study population. Therefore, when using negative control exposures within the unexposed to detect bias in LASSO PS weighted estimators, we suggest evaluating both the bias and the percent bias (bias relative to the bias of the unadjusted estimate) to better understand how the bias within negative control exposure studies relates to the study cohort.
2.4.2 A framework for generating synthetic negative control exposure cohorts
Definition 2 requires negative control exposures to be a function of the true propensity score, $e(X)$, which is rarely known. Instead, we can simulate a synthetic negative control exposure with odds that are proportional to the estimated odds of exposure in the full population. This is stated more formally in the following definition.
Definition 3.
Let $i=1,...,n$ index a sample of $n$ independent and identically distributed random variables with data structure, $\{Y,X,A\}$, where $Y$, $X$, $A$, and $e(X)$ are defined in Section 2.1. Let $\widehat{e}(X)_{n}$ represent a sample estimator for $e(X)$. We define a synthetic negative control exposure with proportional odds as a binary variable that is generated with probability $\widehat{g}(X)_{n}$, where $\frac{\widehat{g}(X)_{n}}{1-\widehat{g}(X)_{n}}\propto\frac{\widehat{e}(X)_{n}}{1-\widehat{e}(X)_{n}}$.
Synthetic negative control exposures are generally applicable because they are simulated using the sample estimator $\widehat{e}(X)_{n}$. However, this also creates a fundamental limitation in that the confounding structure for synthetic negative control exposures can only reflect information captured in $\widehat{e}(X)_{n}$. Still, if $\widehat{e}(X)_{n}$ consistently estimates $e(X)$, we postulate that synthetic negative control exposures satisfy the conditions of Proposition 2 asymptotically and can be used to detect bias caused by improper undersmoothing of LASSO PS weighted estimators.
Several variations for generating synthetic negative control exposures have been proposed [41, 42, 43, 44]. A common thread across the approaches is that they approximate a confounding structure where exchangeability between the synthetic exposure and observed outcome among the unexposed is closely connected to partial exchangeability in the full study population, as described in Proposition 2. Here, we propose additional modifications to previous frameworks for the purpose of validating LASSO PS-weighted analyses. We outline the framework in the following algorithm and provide comments on various aspects of the framework below.
Algorithm 1 Framework for generating synthetic negative control exposure datasets
1:Fit a LASSO model predicting the exposure in the full population. This model is tuned using cross-validation to minimize prediction error. Let $\widehat{e}(X)$ represent the fitted values from this model.
2:Subset the data to the unexposed group where the counterfactual outcome, $Y^{a=0}$, is observed (i.e., individuals where $A=0$). Throughout, let $n_{u}$ represent the number of individuals in the unexposed group and $n$ the number of individuals in the full study cohort.
3:For each unexposed individual, $i=1,...,n_{u}$, use the fitted model in Step 1 to assign a synthetic exposure probability, $\pi_{i}$, so that 1) the odds of $\pi_{i}$ are proportional to the odds of $\widehat{e}(X)$, and 2) the proportion of those assigned synthetic exposure in the unexposed group is equal to the proportion of those assigned exposure in the study cohort (in expectation). More formally, let $\pi_{i}=\frac{exp(c+\theta_{i})}{1+exp(c+\theta_{i})}$, where $\theta_{i}=log\left(\frac{\widehat{e}(X)}{1-\widehat{e}(X)}\right)$, and $c$ is a constant such that $\sum_{i=1}^{n_{u}}\pi_{i}=\left(\frac{n_{u}}{n}\right)n_{u}$.
4:Take $k$ bootstrapped samples from the unexposed group, where each bootstrapped sample is of size $n$ (i.e., the size of the full study cohort).
5:For each sampled unit, $j$ (where $j=1,...,n$), in the $k^{th}$ bootstrapped sample, conduct a single independent Bernoulli trial with probability $\pi_{j}$ to determine whether sampled unit $j$ is assigned to the synthetic exposure group.
6:For each of the $k$ negative control exposure cohorts, apply alternative LASSO PS weighted analyses. For each analysis, calculate the bias and percent bias (bias relative to the bias of the unadjusted estimate), and take the average across the $k$ cohorts to determine the mean synthetic bias and mean percent synthetic bias for each analysis.
Comment 1.
The framework begins by fitting a LASSO model predicting exposure that is tuned using cross-validation. This model does not implement any undersmoothing and, consequently, is not expected to minimize bias in PS weighted causal estimators. However, this is not the objective since this model is not used for constructing causal estimators. The goal here is to fit the most accurate model for the PS function, $e(X)$, to generate synthetic negative control exposures that mimic the causal structure between covariates and exposure. Ertefaie et al [15] explain that while undersmoothing can result in PS weighted estimators that converge at a faster rate to causal parameters, undersmoothing does not improve the rate of convergence to the PS function, $e(X)$, itself. A LASSO model that is tuned using cross-validation will estimate the PS regression more precisely, whereas an undersmoothed LASSO model sacrifices some precision of that estimate in the interests of enhancing precision of estimation of the eventual causal parameter. Therefore, we do not undersmooth the fitted LASSO model in Step 1 to avoid degrading the accuracy of the fitted model in terms of approximating the true PS function, $e(X)$.
Comment 2.
A critial aspect of the framework is how bootstrapping is applied (Step 4). First, it is important that bootstrapping is applied before assigning synthetic negative control exposures (Step 5). This is consistent with a model-based bootstrap for exposure, where exposure is regenerated (simulated) for each bootstrapped sample. Regenerating rather than resampling exposure is necessary to avoid known problems that occur when applying LASSO models to nonparametric bootstrap samples [45, 46]. Regenerating exposure status is also necessary to avoid issues with nonpositivity in plasmode datasets that has been discussed extensively in more recent work [47]. Second, because bootstrapping is restricted to the unexposed group, traditional bootstrap applications ($n$-out-of-$n$ bootstrap) restrict the ability to evaluate the performance of estimators in cohorts where the sample size resembles the study of interest. To address this, the framework uses bootstrap oversampling. While less common, bootstrap oversampling can be used when the objective is to evaluate the performance of estimators in populations that are larger than the cohort from which bootstrap samples are taken [48, 49, 50]. Although bootstrap approaches that result in samples with many repeated observations can have limitations in certain settings [32], we suggest that using a model-based bootstrap for exposure helps to mitigate these limitations for PS estimators. Still, Step 4 in Algorithm 1 could accommodate alternative subsampling techniques when one is concerned with the performance of bootstrap oversampling in more extreme settings [51, 52]. Limitations of bootstrap oversampling and alternative subsampling approaches are discussed further in Section 5.
Comment 3.
Finally, because the synthetic datasets do not approximate the full confounding structure but rather a confounding structure that is related to the counterfactual population under no exposure, the framework is only used for bias detection rather than trying to evaluate a range of statistical properties to select the ‘best’ analytic approach for the study at hand. If an undersmoothed LASSO PS weighted estimator is unable to adequately control for confounding captured by the cross-validated LASSO model from Step 1 to produce unbiased synthetic effect estimates, it is unlikely the same estimator will satisfy partial exchangeability when applied to the original study population.
3 Evaluation
We used a series of Monte Carlo simulation experiments to investigate the performance of alternative approaches for tuning LASSO models for PS weighting, and the use of synthetic negative control exposures to detect biased analyses. We considered two different simulation setups where the exposure effect was simulated to be null. We simulated datasets under the null because the goal of synthetic negative control exposures is to detect bias due to violations of partial exchangeability, which is related to bias in the counterfactual population under no exposure. Limitations of using negative controls when bias in this counterfactual population does not correlate strongly with bias in the study population are discussed in Section 5.
The first simulation setup was motivated from Wyss et al [20] where the data structure was simulated to reflect settings common in healthcare database studies where the prevalence of exposure is much greater than the outcome incidence and where the vast majority of baseline features available for covariate adjustment are spurious binary indicators (sparse high-dimensional data structures). The second simulation setup comes directly from one of the simulations in Ertefaie et al [15]. This second simulation included a smaller set of baseline covariates, but where the distribution of the baseline covariates and their association with exposure were more complex (continuous covariates with nonlinear associations with exposure). We briefly outline each simulation setup below. Details along with R software code are provided in the Supplemental Appendix.
  • Simulation Setup 1: We simulated a binary exposure, $A$, a binary outcome, $Y$, and 1000 baseline covariates, where $e(X)=expit(\beta_{0}+\beta_{1}X_{1}+\beta_{2}X_{2}+\ldots+\beta_{100}X_{100})$ and $P(Y|X)=expit(\alpha_{0}+\alpha_{1}X_{1}+\alpha_{2}X_{2}+\ldots+\alpha_{100}X_{100})$. Each coefficient in the exposure and outcome model was drawn from a separate $unif(0,0.693)$ distribution. Each $X_{k}\sim Bernoulli(0.2)$, where ($k=1,\ldots,100$). We simulated an additional 900 variables binary variables ($X_{101}$ through $X_{1000}$) that had no effect on exposure or outcome (spurious variables). The prevalence of exposure and incidence of the outcome were simulated to be 30% and 5%, respectively. We considered sample sizes of 5,000, 10,000, 20,000, and 40,000.
  • Simulation Setup 2: We simulated a binary exposure, $A$, and a normally distributed outcome, $Y$, and 10 baseline covariates, $X_{1}$ through $X_{5}\sim unif(-2,2)$ and $X_{6}$ through $X_{10}\sim Bernoulli(0.6)$. Exposure was generated with probability $e(X)$ where $e(X)=expit(X_{2}^{2}-exp(0.5X_{1})-X_{3}+X_{4}-exp(0.5X_{5})+X_{6}+X_{7})$. The outcome model was simulated as $Y=-2X_{2}^{2}+2X_{1}+2E(X_{2}^{2})+X_{2}+X_{1}X_{2}+X_{3}+X_{4}+2X_{5}^{2}-2E(X_{5}^{2})+\epsilon$, where $\epsilon\sim\mathcal{N}(0,0.1)$. Covariates $X_{8}-X_{10}$ were spurious (unrelated to exposure and outcome). We considered sample sizes of 1,000, 2,000, 5,000, and 10,000.
For both simulations, we generated 500 datasets. For Simulation 1, we estimated the PS using LASSO regression that included the main effects of all 1000 variables. For Simulation 2, we used the Highly Adaptive Lasso with default tuning parameters from the hal9001 R package [53] to generate an expanded set of indicator basis functions, $W$. We then fit LASSO models with different tuning criteria on the generated features, $W$. Tuning approaches included cross-validation and the previously described balance metrics. For each fitted model, we estimated the mean difference in outcomes across exposure groups using the following PS weighting approaches [54, 55]. Example R code is provided in the Supplemental Appendix.
  • Inverse Probability Weighting: $w_{i}=\frac{A_{i}}{e(X_{i})}+\frac{(1-A_{i})}{(1-e(X_{i}))}$
[…]
  • Overlap Weights: $w_{i}=A_{i}(1-e(X_{i}))+(1-A_{i})e(X_{i})$
Previous work has shown that cross-fitting or using out-of-sample predictions when estimating nuissance functions can improve the properties of effect esitmates [20, 56, 57, 58]. Therefore, for all LASSO models, we used out-of-fold (out-of-sample) predictions when assigning weights.
For each simulated dataset, we then ran Algorithm 1 described previously to produce synthetic negative control exposures for bias detection. For Step 4 in Algorithm 1, we took 500 bootstrapped samples from each simulated dataset to generate 500 synthetic negative control exposure cohorts. We applied analyses to each of the generated synthetic cohorts to calculate the estimated bias and percent bias for that simulated dataset. We then averaged across all the simulated datasets to evaluate how well the synthetic bias and synthetic percent bias aligned with the actual bias and percent bias for the two simulation setups described above.
4 Results
Figures 2 and 3 show the bias (Plot A) and percent bias (Plot C) in effect estimates along with the corresponding synthetic bias (Plot B) and synthetic percent bias (Plot D) averaged across all simulated datasets. For both scenarios, undersmoothing the LASSO models using balance diagnostics resulted in less bias in estimated exposure effects when compared to tuning using cross-validation. Figures 2 and 3 show that as the sample size increased the bias for all undersmoothed PS weighted analyses approached zero at a faster rate compared to the bias in the cross-validated PS weighted analyses (Plots A and C). The performance between the two balance criteria for undersmoothing the LASSO models was similar.
Figures 2 and 3 further show that general patterns in bias between the synthetic negative control studies and study population were similar (both in terms of bias, shown in Plots A and B, and relative or percent bias, shown in Plots C and D). However, while overall patterns in bias were similar, the magnitude of bias within the synthetic negative control studies (synthetic bias) was smaller than the bias in the study population (Plots A and B). As the sample size increased, differences in the magnitude of bias between the synthetic cohorts vs the actual study population decreased.
Figure 2: Bias and relative (percent) bias in effect estimates (Plots A and C) and synthetic negative control effect estimates (Plots B and D) for Simulation Setup 1.
Figure 3: Bias and relative (percent) bias in effect estimates (Plots A and C) and synthetic negative control effect estimates (Plots B and D) for Simulation Setup 2.
The closer alignment in the magnitude of bias within the study population vs the synthetic negative control studies is a result of how closely covariate differences within the synthetic cohorts approximate those in the study population, which is illustrated in Figures 4 and 5. Figure 4 shows differences in all 1000 covariates across exposure groups plotted against differences in the same covariates across the synthetic exposure groups for one simulated dataset. Covariate differences across the synthetic exposure groups were averaged across the 500 generated synthetic cohorts for that dataset. Figure 5 shows differences in the expanded set of indicator basis functions across exposure groups plotted against differences in the same features (indicator basis functions) across the synthetic exposure groups for one simulated dataset. Differences in the indicator basis functions across the synthetic exposure groups were averaged across the 500 generated synthetic cohorts for that dataset.
Both Figures 4 and 5 show that covariate differences (or differences in the indicator basis functions) across the synthetic exposures were more closely aligned with those across the actual exposure groups as the sample size increased. This is illustrated by the coefficient of determination (R-squared) moving closer to 1, and the slope of the least squares regression line (red line in Plots A-D) becoming more closely aligned with a slope of 1 (blue line in Plots A-D representing perfect alignment). This is because as the sample size increased, the LASSO model used to generate the synthetic negative control exposure groups more closely approximated the true PS function (Step 1 in Algorithm 1). This, in turn, resulted in synthetic negative control cohorts where patterns of confounding (covariate differences) more closely reflected those in the actual study cohort.
Still, even for large sample sizes (Plot D in Figures 3 and 4), covariate differences across the synthetic exposure groups did not mirror those across the actual exposure groups, but were only an approximation as expected. As covariate differences across synthetic exposure groups more closely align with differences across exposure groups in the study population, the magnitude of the synthetic bias more closely reflects the magnitude of bias in the study population (Plots A and B in Figures 2 and 3). It is interesting, however, that even for the scenarios where the magnitude of bias between the synthetic cohorts and actual population differed by a large amount (e.g., smallest sample size in Figures 2 and 3), the relative performance between the different estimators, which is best reflected in the percent bias, was still a close approximation to the relative performance of the estimators in the study population.
Figure 4: Covariate differences plotted across exposure groups and synthetic negative control exposure groups for one simulated dataset for Simulation Setup 1. Plots A-D show covariate differences for a dataset that had a sample size of 5K (Plot A), 10K (Plot B), 20K (Plot C), and 40K (Plot D).
Figure 5: Differences in the indicator basis functions that were generated for the Highly Adaptive Lasso plotted across exposure groups and synthetic negative control exposure groups for one simulated dataset for Simulation Setup 2. Plots A-D show covariate differences for a dataset that had a sample size of 1K (Plot A), 2K (Plot B), 5K (Plot C), and 10K (Plot D).
Finally, in Figure 6 we present absolute standardized differences in covariates across exposure groups in the study population before and after PS weighting for Simulation Setup 1. Here, we only show balance plots when using inverse probability weights and a LASSO model that was tuned using cross validation to illustrate how balance diagnostics, by themselves, can be inadequate to inform investigators what level of covariate balance is necessary to adequately remove measured confounding bias for the given study and analytic approach.
Figure 6: Absolute standardized difference of each covariate across exposure groups for Simulation Setup 1. Plots A through D correspond to sample sizes 5K (Plot A), 10K (Plot B), 20K (Plot C), and 40K (Plot D). The x-axis for each plot shows standardized differences before PS weighting (unadjusted), while the y-axis shows the standardized differences after PS weighting. The red horizontal dotted line indicates the value that is commonly used to determine adequate covariate balance after PS weighting (standardized difference $<0.1$). For the plots shown here, weighting was done using inverse probability weights and the Lasso model used to estimate the PS was tuned using cross validation.
Figure 6 shows that for all sample sizes in Simulation Setup 1, the LASSO model that was tuned using cross-validation resulted in standardized differences in covariates that were well below the commonly used threshold of 0.1 after inverse probability weighting [23]. However, as shown previously in Figure 2, the inverse probability weighted estimator with a cross-validated LASSO model also resulted in large bias in estimated exposure effects, particularly for smaller sample sizes. This example illustrates that it can be difficult to know how much balance is necessary to adequately remove confounding bias for a given study. Supplementing balance diagnostics with results from synthetic negative control studies (Figure 2 Plots B and D) can help investigators determine if the level of balance achieved adequately removes measured confounding bias for the given study and analytic approach.
5 Discussion
In this study, we considered using balance criteria to determine the degree of undersmoothing when fitting LASSO models for PS weighting. Because no tuning criteria and weighting approach are univerally best, we further proposed a framework to generate synthetic negative control exposures to detect bias caused by improper undersmoothing for the given study and PS weighting approach. Numerical experiments suggest that using balance criteria to undersmooth LASSO models can reduce bias in PS weighted estimators compared to estimators that are tuned using cross-validation. Numerical studies further suggest that synthetic negative control exposures can be useful for bias detection.
Outcome-blind diagnostics are critical for robust and transparent comparisons of design and analytic choices in causal inference [27, 59]. The use of synthetic negative control exposures for bias detection allows investigators to objectively evaluate and compare alternative LASSO PS-weighted analyses in their ability to control for measured confounding without being influenced by estimated exposure effects in the full study population. Consequently, the framework maintains objectivity in study design by not allowing information on the exposure–outcome association to contribute to decisions on model selection [27].
A few limitations deserve attention. It is important to highlight that synthetic negative controls are limited in that they can only detect violations of partial exchangeability caused by lack of control for measured confounders (which can result from improper undersmoothing). Partial exchangeability is necessary for identificaiton of average causal effects, but it is not necessarily sufficient for identification of causal effects. If bias is not detected within synthetic negative control exposure studies, it cannot be determined that the analyses are necessarily valid. Consequently, the framework is only useful as a bias detection or screening tool. Still, this is analogous to how real negative control studies are typically used [60]. The difference, of course, being that synthetically generated datasets can only screen for bias captured in the models used to generate the synthetic data (e.g., measured confounding).
Finally, it is important to highlight that future work could explore variations of the proposed framework for generating synthetic negative control exposures. In particular, the proposed algorithm can be flexible in terms of what sampling technique is applied. Here, we proposed the use of bootstrap oversampling to generate samples of the same size as the original study cohort on which to evaluate estimators. However, if one is concerned about evaluating estimators in samples that contain many repeated observations, future work could consider alternative subsampling techniques, such as the $m$-out-of-$n$ bootstrap or $m$-out-of-$n$ subsampling without replacement [51, 52]. Subsampling techniques are widely regarded as being more robust than the traditional $n$-out-of-$n$ bootstrap (and bootstrap oversampling) as they are asymptotically valid under weaker conditions [51, 52, 32]. But subsampling techniques require one to select the size of the subsamples which involves some exercise of judgment and are limited in that the size of the subsamples must be smaller than the original study cohort. Future work could also consider variations that re-estimate the CV LASSO used to generate synthetic exposure probabilities within each bootstrapped sample.
In summary, both theory and simulations have shown that undersmoothing LASSO models can reduce bias of PS weighted estimators. We conclude that the use of balance diagnostics to determine the degree of undersmoothing when fitting LASSO PS models, and the use of synthetic negative control exposures to detect bias caused by improper undersmoothing are promising tools to improve confounding control for large-scale PS weighted analyses.
Acknowledgements
This work was funded by National Institutes of Health grant NIH RO1LM013204 and Patient-Centered Outcomes Research Institute contract PCORI ME-2022C1-25646.
S1   Supplemental Appendix
S1.1   General Overview of the Highly Adaptive Lasso
HAL is a machine learning prediction algorithm that can be used to nonparametrically estimate regression functions. To understand HAL, it is necessary to understand the conditions (or assumption) required by HAL. First, HAL requires the total variation of a function (or variation norm) to be bounded, where the total variation can be thought of as a measure of the complexity of a function. For example, in the simple case of a monotone function, $f(\cdot)$, on the interval $[0,1]$, the total variation is simply $|f(1)-f(0)|$. For more general functions, the total variation is equal to the cumulative sum of all the absolute incremental changes in the function’s values over its domain. Intuitively, if we plotted a non-linear/non-monotone continuous univariate function and traced the function with a piece of string, the total variation of the function can be thought of as the length of the string. Being able to bound the total variation of a function ensures that the function cannot wiggle around too much. The concept of bounded total variation can be extended to higher dimensions (and non-continuous functions), but the general idea remains the same.
By requiring the total variation of the function to be bounded, HAL places a global constraint on the behavior of the function rather than local constraints. This distinction is important to understanding HAL. The former controls how much the function can fluctuate globally, while the latter only controls how much the function can fluctuate locally at each point on its domain. One can globally constrain a function by imposing very strong local smoothing constraints; for example, by requiring the function to be many times differentiable with bounded derivatives. However, as the dimension grows, the smoothness constraints needed to globally constrain the function become very strong. This leads to the curse of dimensionality where conventional methods fail to estimate high dimensional functions at a fast enough rate. The main challenge for local smoothing methods is that they try to impose a global constraint on the function by imposing increasingly restrictive local constraints. HAL circumvents this issue by imposing a global constraint directly without imposing any local smoothness constraints.
In addition to requiring the total variation to be bounded, HAL requires some regularity conditions on the function. Mainly, that the function is cadlag, meaning that it is mostly continuous everywhere but can jump finitely many times. Cadlag functions are very general and do not require local smoothness (allow discontinuities). Understanding this function class is also key to understanding HAL since the implementation of HAL is based on recognizing two features of cadlag functions of bounded total variation. First, they can be approximated arbitrarily well by linear combinations of indicator jump functions (e.g., $\mathbbm{1}(X\geq x)$). Second, the total variation (or variation norm) of a linear combination of indicator jump functions is equal to the absolute sum of the regression coefficients in front of the indicator jump functions.
In the context of a propensity score function, $e(X)$, that is cadlag with a bounded variation norm, Benkeser & van der Laan (2016) show that the baseline covariates, $X$, can be expanded into a series of $n(2^{d}-1)$ binary indicator variables (i.e., indicator basis functions), $W$, such that as $n\rightarrow\infty$ the logit of the propensity score function, $logit(e(X))$, can be approximated arbitrarily well by a linear combination of the binary indicators written as:
$logit(g(W))=\gamma_{0}+W^{\top}\gamma$ (4)
where $g(W)=P(A|W)$, $\gamma$ is a $n(2^{d}-1)$ dimensional vector of parameters, and $\gamma_{0}$ is a scalar. For a theoretical explanation on the construction of the indicator basis functions, $W$, see Benkeser & van der Laan (2016) and Ertefaie et al (2022).
In real world settings with finite sample size, it is not possible to adjust for the entire set of binary features in $W$. Consequently, some dimension reduction is needed to approximate $g(W)$. Benkeser and van der Laan (2016) show that LASSO regression serves this purpose through regularization and provides theoretical guarantees on fast convergence rates. As a result, HAL simply defines the optimization problem as a Lasso regression over the transformed binary indicators, $W$. In the context of estimating the PS, HAL becomes an L1-regularized logistic regression, where the parameter vector, $\gamma$, in Equation 4 is estimated based on the following penalized likelihood:
${\cal L}(\gamma)=\sum_{i=1}^{n}-A_{i}log(p(A_{i}|W_{i});\gamma)-(1-A_{i})log(1-p(A_{i}|W_{i};\gamma))+\lambda\sum_{j=1}^{m}\mid\gamma_{j}\mid$ (5)
where $n$ is the sample size and $m$ is the number of parameters in $\gamma$, which can be up to $n(2^{d}-1)$. As $n\rightarrow\infty$, HAL will converge to the logistic model defined in Equation 4, but in finite samples HAL is just an approximation to $g(W)$.
S1.2   Proof of Proposition 1
Proposition 1: Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure. If $Y^{a}\perp\!\!\!\perp A|X$ it follows that $Y\perp\!\!\!\perp Z|X$.
Proof.
Let $Y^{a}\perp\!\!\!\perp A|X$. From the definition of a negative control exposure, $P(Z|X)$ can be written as a 1-to-1 function of $P(A|X)$, and there is no unmeasured common cause of $Z$ and $Y$. This implies that $X$ is sufficient to satisfy $Y^{z}\perp\!\!\!\perp Z|X$. From the definition of a negative control exposure, we further know that $Y^{(a,z)}=Y^{a}$ (no causal effect of $Z$ on $Y$). This further implies that $Y^{z=0}=Y^{z=1}=Y$. Therefore, $Y\perp\!\!\!\perp Z|X$
S1.3   Proof of Proposition 2
Proposition 2: Assume we have the data structure, $\{Y,X,A,Z\}$, where $Z$ is a negative control exposure. If $Y^{a=0}\perp\!\!\!\perp A|X$ it follows that $Y\perp\!\!\!\perp Z|X,A=0$.
Proof.
Let $Y^{a=0}\perp\!\!\!\perp A|X$. From the definition of a negative control exposure, $Y^{a}\perp\!\!\!\perp Z|X$ and $A\perp\!\!\!\perp Z|X$. This implies that $Y^{a=0}\perp\!\!\!\perp Z|X,A=0$. Therefore, $Y\perp\!\!\!\perp Z|X,A=0$ by consistency. ∎
S1.4   Proof for Comment 2 on Definition 2 in Section 2.4.1
Assume we have the data structure defined above, where $Y^{a=0}\perp\!\!\!\perp A|X$ and assume the relationship between $X$ and $A$ is linear (or linear on the log-odds scale). Let $D_{1}=E[X|A=1]-E[X|A=0]$ represent the mean difference in the covariates, $X$, across exposure groups and let $D_{2}=E[X|Z=1,A=0]-E[X|Z=0,A=0]$ represent the mean difference in $X$ across synthetic exposure groups after restricting on $A=0$. Further, let $B_{1}=E[Y^{a=0}|A=1]-E[Y^{a=0}|A=0]$ represent the bias caused by $X$ on the effect of $A$ on $Y^{a=0}$ on the risk difference scale (bias in the unadjusted risk difference), and let $B_{2}=E[Y|Z=1,A=0]-E[Y|Z=0,A=0]$ represent the bias caused by $X$ on the effect of $Z$ on $Y$ within the unexposed population on the risk difference scale (bias in the synthetic unadjusted risk difference). If $D_{1}=D_{2}$ it follows that $B_{1}=B_{2}$.
Proof.
Let $Y^{a=0}\perp\!\!\!\perp A|X$. Using the generalized bias formulas for uncontrolled confounding developed by Vanderweele & Arah (2011), the magnitude of bias caused by $X$ on the effect of $A$ on $Y^{a=0}$ on the risk difference scale can be expressed as
$\sum_{x}\{P[Y^{a=0}|A=a,X=x]-P[Y^{a=0}|A=a,X=x^{\prime}]\}\{P(X|A=1)-P(X|A=0)\}P(X=x)$
If we make the simplifying assumption that the relationship between $X$ and $A$ is linear (or linear on the log-odds scale), the this implies that $P(X|A=1)-P(X|A=0)$ does not vary between strata of $X$. Under this condition, Vanderweele & Arah (2011) show that the above expression then simpifies to
$\{P[Y^{a=0}|A=a,X=x]-P[Y^{a=0}|A=a,X=x^{\prime}]\}\{P(X|A=1)-P(X|A=0)\}$
Linearity between $X$ and $A$ also implies that the above expression further reduces to
$\{P[Y^{a=0}|A=a,X=x]-P[Y^{a=0}|A=a,X=x^{\prime}]\}\{E(X|A=1)-E(X|A=0)\}$
Similarly, the bias caused by $X$ on $Z$ and $Y$ among the unexposed (i.e., those with $A=0$) can be expressed as
$\{P[Y|X=x,A=0]-P[Y|X=x^{\prime},A=0]\}\{E(X|Z=1,A=0)-E(X|Z=0,A=0)\}$
The last two expressions are equal since $P[Y|X=x,A=0]=P[Y^{a=0}|X=x,A=0]=P[Y^{a=0}|X=x]$ and we assume that $E(X|Z=1,A=0)-E(X|Z=0,A=0)=E(X|A=1)-E(X|A=0)$
S1.5   Example R Code
##    Ψdat_gen: function to generate data
### scenario = 1, high-dimensional setting
### scenario = 2, observational study with 10 covariates for HAL implenetation (Ertefaie et al. 2023)
dat_gen <- function(n, ps, seed1, seed2){
   if(scenario == 1){### high-dimensional data
    ΨΨnstudy<- n
    ΨΨnvars<- 1000
    ΨΨ## defining alpha and beta coefficients
    ΨΨnc<- 100
    ΨΨni<- 0
    ΨΨnr<- 0
    ΨΨns<- nvars-(nc+ni+nr)
    ΨΨ## global parameters for sim (only want to set once so they are same for all sims)
    ΨΨ## seed2 should remain same across simulation runs
    ΨΨset.seed(seed2)
    ΨΨcoef_strength<- 0.693
    ΨΨalpha_conf<- runif(nc, 0.0, coef_strength)
    ΨΨbeta_conf<- runif(nc, 0.0, coef_strength)
    ΨΨalpha_temp<- c(alpha_conf)
    ΨΨbeta_temp<- c(  beta_conf)
    ΨΨrandom_neg<- sample(1:length(alpha_temp), 0.5*length(alpha_temp), replace=FALSE)
    ΨΨalpha_temp[random_neg]<- -1*alpha_temp[random_neg]
    ΨΨbeta_temp[random_neg]<-  -1*beta_temp[random_neg]
    ΨΨalpha<-  matrix(c(alpha_temp, rep(0, ns)), ncol=1)
    ΨΨbeta<-   matrix(c(beta_temp, rep(0, ns)), ncol=1)
    ΨΨbetaE<- 0.0
    ΨΨcprev<- runif(nvars, 0.01, 0.1)
    ΨΨoprev<- 0.05
    ΨΨtprev<- 0.30
    ΨΨ## resetting seed so that it is unique for each simulation
    ΨΨ## seed1 should change each run to get new data
    ΨΨset.seed(seed1)
    ΨΨ# generate synthetic matrix of baseline covariates
    ΨΨXcovs<- matrix(NA, nrow=nstudy, ncol=nvars)
    ΨΨfor(pp in 1:nvars){
    ΨΨΨXcovs[,pp]<- rbinom(nstudy, 1, cprev[pp])
    ΨΨ}
    ΨΨXcovs<- as.data.frame(Xcovs)
    ΨΨnames(Xcovs)<- c(paste0(’x’, 1:nvars))
    ΨΨW<- as.matrix(Xcovs)
    ΨΨlinear_pred_e<- W %*alpha
    ΨΨlinear_pred_y<- W %*beta
    ΨΨ##function to find intercept to get specified treatment prevalence
    ΨΨtreatment_inc<- tprev
    ΨΨfn <- function(c) mean(plogis(c + linear_pred_e)) - treatment_inc
    ΨΨalpha0 <- uniroot(fn, lower = -20, upper = 20)\$root
    ΨΨEe <- (1 + exp( -(alpha0 + linear_pred_e) ))^-1
    ΨΨe<- rbinom(nstudy, 1, Ee)
    ΨΨA<- e
    ΨΨp<- Ee
    ΨΨ##function to find intercept to get specified outcome incidence
    ΨΨoutcome_inc<- oprev
    ΨΨfn <- function(c) mean(plogis(c + betaE*e + linear_pred_y  )) - outcome_inc
    ΨΨbeta0 <- uniroot(fn, lower = -20, upper = 20)\$root
    ΨΨEy <- (1 + exp( -( beta0 + betaE*e + linear_pred_y )))^-1
    ΨΨY <- rbinom(nstudy, 1, Ey)
    ΨΨY1<- (1 + exp( -( beta0 + betaE*1 + linear_pred_y )))^-1
    ΨΨY0<- (1 + exp( -( beta0 + betaE*0 + linear_pred_y )))^-1
  Ψ}
 Ψif(scenario == 2){### Simulation in Supplemental Material of Ertefaie et al. (2023)
    ΨΨset.seed(seed1)
    ΨΨx1 = runif(n,-2,2)
    ΨΨx2 = runif(n,-2,2)
    ΨΨx3 = runif(n,-2,2)
    ΨΨx4 = runif(n,-2,2)
    ΨΨx5 = runif(n,-2,2)
    ΨΨx6 = rbinom(n,1,0.6)
    ΨΨx7 = rbinom(n,1,0.6)
    ΨΨx8 = rbinom(n,1,0.6)
    ΨΨx9 = rbinom(n,1,0.6)
    ΨΨx10 = rbinom(n,1,0.6)
    ΨΨ##function to find intercept to get specified treatment prevalence
    ΨΨlinear_pred_e = (1*x2^2-exp(x1/2)-x3+x4-exp(x5/2)+x6+x7)/2
    ΨΨtprev<- 0.50
    ΨΨtreatment_inc<- tprev
    ΨΨfn <- function(c) mean(plogis(c + linear_pred_e)) - treatment_inc
    ΨΨalpha0 <- 0 #uniroot(fn, lower = -20, upper = 20)\$root
    ΨΨp <- (1 + exp( -(alpha0 + linear_pred_e) ))^-1
    ΨΨA = 1-rbinom(n, size = 1, prob = p)
    ΨΨY_error<- rnorm(n, mean = 0, sd = 0.1)
    ΨΨY1 = +0*1 +(-2*x2^2+2*x1+2*mean(x2^2)+x2+x1*x2+x3+x4+2*x5^2-2*mean(x5^2))/1 + Y_error
    ΨΨY0 = +0*0 +(-2*x2^2+2*x1+2*mean(x2^2)+x2+x1*x2+x3+x4+2*x5^2-2*mean(x5^2))/1 + Y_error
    ΨΨY  = A*Y1 + (1-A)*Y0
  Ψ}
   if(ps == 1) {return(data.frame(Y, Y1, Y0, A, p, Xcovs))}
   if(ps == 2) {return(data.frame(Y, Y1, Y0, A, p, x1, x2, x3, x4, x5, x6, x7, x8, x9, x10))}
##    Ψtreatment_model: function to to fit lasso model for many values of lambda
#’  Calls glmnet to fit propensity score models corresponding to different degrees of regularization
#’ @param data a dataset or matrix containing baseline covariates
#’ @param treatment a binary vector for treatment
#’ @param foldid fold each subject belongs to
#’ @param alpha the elasticnet tuning parameter defined in glmnet (default is 1 for Lasso)
#’ @param lambda_ratio the ratio from the largest to smallest lambda to consider (constrains the range of lambda values)
#’ @param nlambda the number of lambda tuning parameters to consider when fitting glmnet
#’ @param maxit maximum number of iterations as defined in glmnet
#’ @param nmodels the number of undersmoothed models to return PS values for
#’ @param penalty optional vector to specify different penalties for variables like in adaptive lasso (default is NULL)
#’ @param par optional TRUE/FALSE to implement parallel computing (default is FALSE)
treatment_model<- function(data,
                           treatment,
                           foldid,
                           alpha=1,
                           lambda_ratio=NULL,
                           nlambda=100,
                           nmodels=NULL,
                           maxit=5000,
                           penalty=NULL,
                           par=FALSE){
     Wmat = as.matrix(data)
     sx = Matrix::Matrix(Wmat, sparse=TRUE)
     if(is.null(penalty)){
          penalty_factor<- rep(1, ncol(Wmat))
     if(!is.null(penalty)){
          penalty_factor<- penalty
     if(is.null(lambda_ratio)){
          lambda_ratio<- ifelse(nrow(sx) < ncol(sx), 0.01, 1e-04)
     glmnet.e<- NULLΨΨ
     glmnet.e<-  glmnet::cv.glmnet(x = sx,
                                   y = treatment,
                                   family = "binomial",
                                   type.measure = "deviance",
                                   alpha = alpha,   ##ridge regression alpha=0, lasso alpha=1
                                   nlambda = nlambda,
                                   lambda.min.ratio = lambda_ratio, #ifelse(nrow(sx) < ncol(sx), 0.01, 1e-04),
                                   #nfolds = 10, ##don’t specify this when using foldid
                                   #penalty.factor = c(rep(1, dim(sx)[2])),
                                   penalty.factor = penalty_factor,
                                   parallel = par,
                                   standardize = FALSE,  ###HAL does not standardize design matrix
                                   maxit=maxit, #5000,
                                   foldid = foldid,
                                   keep=TRUE)
Ψ    ## extracting predicted values & lambda starting and stopping values
Ψ    gns1<- NULL
Ψ    gns2<- NULL
Ψ    gns1<- predict(glmnet.e, newx = Wmat, s=glmnet.e\$lambda, type = "response")  ##predicted values for each lambda
Ψ    gns2<- plogis(glmnet.e\$fit.preval[, 1:ncol(gns1)]) ## Cross Validated (out-of-fold) predicted values for each lambda
Ψ    ## lambda value that optimizes CV prediction (minimizes CV prediction error)
Ψ    n.lambda.start<- which(glmnet.e\$lambda == glmnet.e\$lambda.min)
Ψ    lambda.start<- glmnet.e\$lambda.min
Ψ    ## only keeping lambda values that are less than or equal to lambda that optimizes CV prediction
Ψ    preds_under1<- gns1[,n.lambda.start:ncol(gns1)]  ##same-sample predicted values
Ψ    preds_under2<- gns2[,n.lambda.start:ncol(gns2)]  ##out-of-fold predicted values
Ψ    lambda_vector<- glmnet.e\$lambda[n.lambda.start:length(glmnet.e\$lambda)]
Ψ    lassocoef = glmnet.e\$glmnet.fit\$beta[,n.lambda.start:length(glmnet.e\$lambda)]
Ψ    coef_mat<- lassocoef
 Ψ    ##number of selected variables for each lambda value
Ψ    n_selected_vars<- apply(lassocoef, 2, function(x){sum(x != 0)})
Ψ    results<- list(preds_under1,
                    preds_under2,
                    coef_mat,
                    lambda_vector,
                    n_selected_vars)
  Ψ  names(results)<- c(’preds’,
                        ’preds_cf’,
                        ’coef_mat’,
                        ’lambdas’,
                        ’n_vars’)
  Ψ  return(results)
##     Functions to calculate covariate balance:
##           weighted.var() and mw_fun() are helper functions used within weighted_diff()
##           weighted_diff() is a helper function used within balance_weighted_diff()
##           balance_weighted_diff() calculates covariate balance for each covariate
#Helper function used within weighted.diff
#weighted.var function from Gavin Simpson. URL: https://stat.ethz.ch/pipermail/r-help/2008-July/168762.html
weighted.var <- function(x, w, na.rm = FALSE) {
  if (na.rm) {
    w <- w[i <- !is.na(x)]
    x <- x[i]
  sum.w <- sum(w)
  sum.w2 <- sum(w^2)
  mean.w <- sum(x * w) / sum(w)
  (sum.w / (sum.w^2 - sum.w2)) * sum(w * (x - mean.w)^2, na.rm = na.rm)
#Helper function used within weighted.diff
mw_fun<- function(score, treatment){
  score_data<- as.data.frame(cbind(score, (1-score)))
  numerator<- apply(score_data, 1, min)
  weight<- numerator / (treatment*score + (1-treatment)*(1-score))
  return(weight)
#’  Helper function used within ’balance_weighted_diff()’ to calculate weighted standardized differences
#’ @param data a dataset or matrix containing baseline covariates
#’ @param data0 a dataset or matrix containing baseline covariates for unexposed group
#’ @param data1 a dataset or matrix containing baseline covariates for exposed group
#’ @param score a dataset or matrix of fitted propensity score values (each column corresponds to different model)
#’ @param treatment a vector of binary indicators indicating treatment status
#’ @param method weighting method used to calculate weighted standardized differences
#’ @param normalized boolean TRUE/FALSE to indicate use of normalized weights (default is TRUE)
weighted_diff<- function(data,
                         data0,
                         data1,
                         score,
                         treatment,
                         method,
                         normalized,
                         standardize){
  Ψ    ## IPTW weights
  Ψ    if(method==’iptw’){
    Ψ    Ψe1_mean<- mean(treatment)
    Ψ    Ψe0_mean<- 1-e1_mean
    Ψ    Ψ#weight<- (treatment*e1_mean)/score + ((1-treatment)*e0_mean)/(1-score) #stabilized weight
    Ψ    Ψweight<- treatment/score + (1-treatment)/(1-score)  #unstabilized weight
  Ψ    }
  Ψ    ## matching weights
  Ψ    if(method==’mw’){
    Ψ    Ψweight<- mw_fun(score=score, treatment=treatment)
  Ψ    }
  Ψ    ## overlap weights
  Ψ    if(method==’ow’){
    Ψ    Ψweight<- treatment*(1-score) + (1-treatment)*score
  Ψ    }
  Ψ    ## creating weighted cohorts by treatment group
  Ψ    weight0<- weight[treatment==0]
  Ψ    weight1<- weight[treatment==1]
  Ψ    #############################################################
  Ψ    ##  normalized weighted average (weighted.mean normalizes)
  Ψ    if(normalized==TRUE){
    Ψ    Ψfun0<- function(x){weighted.mean(x,weight0)}
    Ψ    Ψfun1<- function(x){weighted.mean(x,weight1)}
    Ψ    Ψfun0.sd<- function(x){sqrt(weighted.var(x, weight0))}
    Ψ    Ψfun1.sd<- function(x){sqrt(weighted.var(x, weight1))}
    Ψ    Ψ#fun0.sd<- function(x){sqrt(wtd.var(x, weight0, na.rm=TRUE))}
    Ψ    Ψ#fun1.sd<- function(x){sqrt(wtd.var(x, weight1, na.rm=TRUE))}
    Ψ    Ψmean0.w<- apply(data0, 2, fun0)
    Ψ    Ψmean1.w<- apply(data1, 2, fun1)
    Ψ    Ψ## same as above but calculated manually
    Ψ    Ψ#mean0.w<- apply(data0, 2, function(x) sum(x*weight0)/sum(weight0))
    Ψ    Ψ#mean1.w<- apply(data1, 2, function(x) sum(x*weight1)/sum(weight1))
   Ψ    Ψ sd0.w<- apply(data0, 2, fun0.sd)
    Ψ    Ψsd1.w<- apply(data1, 2, fun1.sd)
  Ψ    }
  Ψ    ############################################################################
  Ψ    ## unnormalized weighted balance (Austin does not recommend this approach)
  Ψ    if(normalized != TRUE){
    Ψ    Ψmean0.w<- apply(data0, 2, function(x) mean(x*weight0))
    Ψ    Ψmean1.w<- apply(data1, 2, function(x) mean(x*weight1))
    Ψ    Ψsd0.w<- apply(data0, 2, fun0.sd)
    Ψ    Ψsd1.w<- apply(data1, 2, fun1.sd)
    Ψ    Ψ#sd0.w<- apply(data1, 2, function(x) sd(x*weight0))
    Ψ    Ψ#sd1.w<- apply(data0, 2, function(x) sd(x*weight1))
  Ψ    }
  Ψ    #######################################
  Ψ    ##          Calculating mean difference
  Ψ    if(standardize==TRUE){diff_weight<- (mean0.w-mean1.w) / sqrt((sd0.w^2 + sd1.w^2)/2)}
  Ψ    if(standardize==FALSE){diff_weight<- (mean0.w-mean1.w)}
  Ψ    abs_diff_weight<- abs(diff_weight)
  Ψ    return(diff_weight)
#’  Calculates unadjusted and weighted standardized differences for each covariate
#’ @param data a dataset or matrix containing baseline covariates
#’ @param treatment a vector of binary indicators indicating treatment status
#’ @param ps_dat a datset or matrix of fitted propensity score values (each column corresponds to a different model)
#’ @param method weighting method used to calculate weighted standardized differences
#’ @param normalized boolean TRUE/FALSE to indicate use of normalized weights (default is TRUE)
balance_weighted_diff<- function(data,
                                 treatment,
                                 ps_dat,
                                 method,
                                 normalized,
                                 standardize){
  Ψ    treatment=treatment
  Ψ    data=data
  Ψ    ps_dat=as.data.frame(ps_dat)
  Ψ    method=method
  Ψ    ###############################
  Ψ    ##   Unadjusted balance
  Ψ    data0<-  data[treatment==0,]
  Ψ    data1<-  data[treatment==1,]
  Ψ    data_test0<- data0
  Ψ    data_test1<- data1
  Ψ    mean0<- apply(data_test0, 2, mean)
  Ψ    mean1<- apply(data_test1, 2, mean)
  Ψ    sd0<- apply(data_test0, 2, sd)
  Ψ    sd1<- apply(data_test1, 2, sd)
  Ψ    sdf<- apply(data, 2, sd)
  Ψ    if(standardize==TRUE){diff_crude<- (mean0-mean1) / sqrt((sd0^2 + sd1^2)/2)}
  Ψ    if(standardize==FALSE){diff_crude<- (mean0-mean1)}
  Ψ    abs_diff_crude<- abs(diff_crude)
  Ψ    #####################################
  Ψ    ##   PS weighted balance
  Ψ    bal_avg<- NULL
  Ψ    bal_max<- NULL
  Ψ    dat_diff_weight<- NULL
  Ψ    for(iii in 1:ncol(ps_dat)){
    Ψ    Ψps_select<- ps_dat[,iii]
    Ψ    Ψ## weighted balance (weighted_diff is defined above)
    Ψ    Ψweighted_differences<- weighted_diff(data=data,
                                               data0=data0,
                                               data1=data1,
                                               score=ps_select,
                                               treatment=treatment,
                                               method=method,
                                               normalized=normalized,
                                               standardize=standardize)
          diff_weight<- weighted_differences
          dat_diff_weight<- cbind(dat_diff_weight, diff_weight)
  Ψ    dat_diff_weight2<- cbind(diff_crude, dat_diff_weight)
  Ψ    return(dat_diff_weight2)
##    Ψps_undersmooth_bal: function to choose lambda value based on minimizing balance criteria
#’ Undersmoothing Lasso PS Models using balance diagnostics
#’ @param data a dataset or matrix containing baseline covariates
#’ @param treatment a binary vector for treatment
#’ @param ps_dat a matrix of fitted propensity scores
#’ @param normalized a boolean TRUE/FALSE to use normalized weighting
#’ @param standardize a boolean TRUE/FALSE to use standardized differences
ps_undersmooth_bal<- function(data,
                              treatment,
                              ps_dat,
                              method,
                              normalized=TRUE,
                              standardize=TRUE){
  Ψ## note: balance_select calculates balance with normalized weighted averages
  Ψcov_diff<- balance_weighted_diff(data=data,
                                    treatment=treatment,
                                    ps_dat=ps_dat,
                                    method=method,
                                    normalized=normalized,
                                    standardize=standardize)
  Ψ## exclude first column which is unadjusted (crude) differences
  Ψcov_diff<- cov_diff[,-1]
  Ψ## standardized absolute differences
  Ψcov_diff_abs<- apply(cov_diff, 2, abs)
  Ψ## balance metric 1: minimum ASAMD
  Ψcov_diff_abs_avg<- apply(cov_diff_abs, 2, mean)
  Ψbal_m1_index<- which.min(cov_diff_abs_avg)
  Ψbal_m1_value<- cov_diff_abs_avg[bal_m1_index]
  Ψ## balance metric 2: model with the smallest max standardized difference
  Ψcov_diff_max<- apply(cov_diff_abs, 2, max)
  Ψbal_m2_index<- which.min(cov_diff_max)
  Ψbal_m2_value<- cov_diff_max[bal_m2_index]
  Ψselect_value<- c(bal_m1_value, bal_m2_value)
  Ψselect_index<- c(bal_m1_index, bal_m2_index)
  Ψselect_preds<- ps_dat[,select_index]
  Ψnames(select_preds)<- c(’model1’, ’model2’)
  Ψnames(select_value)<- c(’asamd’, ’max_diff’)
  Ψnames(select_index)<- c(’index1’, ’index2’)
  Ψresults<- list(select_preds,
                  select_value,
                  select_index)
  Ψnames(results)<- c(’predictions’, ’balance’, ’index’)
  Ψreturn(results)
##  helper functions to calculate prediciton diagnostics: NLL and Cstat
# function to calculated negative log-likelihood
nloglik <- function(y, pred, trunc = 0.001) {
  pred <- pmin(pmax(pred, trunc), 1-trunc)
  if (all(y == round(y))) {
    - mean(ifelse(y==1, log(pred), log(1-pred)))
  } else {
    - mean(y * log(pred) + (1-y) * log(1-pred))
# function to calculated auc or c-stat
auc <- function(y, pred) {
  require("ROCR")
  performance(prediction(pred, y), "auc")@y.values[[1]]
##   helper function to create folds stratified by variable: created by Susan Gruber (used in Wyss et al. 2024)
stratifyCVFoldsByYandID <- function (V, Y, id = NULL) {
  Ψ# 1. distribute the ids that have Y = 1 in any of the rows equally among all the folds,
  Ψ# 2. separately, distribute the ids that have Y = 0 for all rows equally among the folds
  Ψ# ensure that V is less than the number of cases
  Ψif (is.null(id)) id <- 1:length(Y)
  Ψcase_status_by_id <- by(Y, id, sum)  # this gives n.unique results, sorted by id #
  Ψcase_ids <- names(case_status_by_id)[ case_status_by_id > 0]
  Ψnoncase_ids <- names(case_status_by_id)[ case_status_by_id == 0]
  Ψif (V > min(length(case_ids), length(noncase_ids))) {
  Ψ  Ψstop("number of observations in minority class is less than the number of folds")
Ψ  }
  ΨvalSet.case_ids <- split(sample(case_ids), rep(1:V, length = length(case_ids)))
  ΨvalSet.noncase_ids <- split(sample(noncase_ids), rep(1:V, length = length(noncase_ids)))
  ΨvalidRows <- vector("list", length = V)
  Ψnames(validRows) <- paste(seq(V))
  Ψfold_id <- rep(NA, length(Y))
  Ψfor (v in seq(V)){
    ΨΨvalidRows[[v]] <- which(as.character(id) %inc(valSet.case_ids[[v]], valSet.noncase_ids[[v]]))
    ΨΨfold_id[validRows[[v]]] <- v
 Ψ }
  Ψreturn(list(validRows = validRows, fold_id = fold_id))
##    Function to fit HAL to generate matrix of indicator basis functions (used for Scenario 2)
# see hal9001 package in R for details
hal_model<- function(X, Y, max_degree, num_knots, nfolds, foldid){
     # fitting HAL
     mod_full<- fit_hal(X=X,
                        Y=Y,
                        X_unpenalized = NULL,
                        max_degree = max_degree,
                        smoothness_orders = 0,
                        num_knots = num_knots,
                        reduce_basis = 0.01,
                        family = c("binomial"),
                        lambda = 10,
                        id = NULL,
                        offset = NULL,
                        fit_control = list(cv_select = FALSE, nfolds = nfolds, foldid = foldid, use_min = TRUE,
                                        lambda.min.ratio = .001, prediction_bounds = "default"),
                        basis_list = NULL,
                        return_lasso = TRUE,
                        return_x_basis = TRUE,
                        yolo = FALSE)
     ## design matrix
     x_basis_full<- mod_full\$x_basis[,-1] ## first column is intercept term (all ones)
     return(x_basis_full)
##ΨΨΨ       Main: Running Simulation by Calling Helper Functions Above and Running Analysis
##ΨΨ(note: code below is just example for one run. Need to run in ’for’ loop (for i in 1:nsim) to get multiple runs
##         could also use parallel processing for multpile runs with minor edits)
library(hal9001)
library(Matrix)
library(glmnet)
library(dplyr)
## ΨGenerate data
scenario<- 1
seed1<- i         ## seed for simulating random data (should be different for each run)
seed2<- 110101    ## seed for setting global parameters for scenario 1 (should stay the same across runs)
dat<- dat_gen(n=5000, ps=scenario, seed1=seed1, seed2=seed2)
##   Creating Folds
nfolds = 10
cvfolds <- stratifyCVFoldsByYandID(V=nfolds, Y = dat\$A)
folds <- cvfolds\$validRows
foldid <- cvfolds\$fold_id
##  Getting baseline covariates (for Scenario 2, requires fitting HAL on full data to extract design matrix)
x_names<- names(dat)[which(substr(names(dat), 1, 1)==’x’)]
X_dat<- dat[,names(dat) %inx_names]
if(scenario == 1){ x_basis_full<- X_dat }
if(scenario == 2){ x_basis_full<- hal_model(X=X_dat, Y=dat\$A, max_degree=2, nfolds=NULL, foldid=foldid, num_knots=c(100, 25)) }
## reducing dimension of basis matrix (cleaning based on prevalence)
temp1<- apply(x_basis_full, 2, mean)
temp2<- (temp1 >= .01 & temp1 <= .99)
x_basis_full<- x_basis_full[,temp2]
##  Fitting glmnet on x_basis from above and getting predicted values for multiple lambda values
lasso_object<- treatment_model(data=x_basis_full,
                               treatment=dat\$A,
                               foldid=foldid,
                               alpha=1,
                               lambda_ratio=0.01, #ifelse(nrow(x_basis_full) < ncol(x_basis_full), 0.01, 1e-04)
                               nlambda=200,
                               nmodels=NULL,
                               maxit=5000,
                               penalty=NULL,
                               par=FALSE)
preds<- NULL
preds_cv<- NULL
preds_all<- NULL
coef_mat<- NULL
lambdas<- NULL
nvars<- NULL
preds<- lasso_object[[1]]       #same-sample predictions
preds_cv<- lasso_object[[2]]    #cross-validated (out-of-fold) predictions
coef_mat<- lasso_object[[3]]    #coefficient matrix (coefficients for each lambda)
lambdas<- lasso_object[[4]]     #lambda values
nvars<- lasso_object[[5]]       #number of variables selected by each lambda
#index values for subset of predictions and lambdas (select a range to reduce computation time instead of selecting all values)
steps<- NULL
steps<- floor(quantile(1:ncol(preds), seq(0, 1, .02)))
preds_all<- preds_cv[,steps]
coef_mat<- coef_mat[,steps]
lambdas<- lambdas[steps]
nvars<- nvars[steps]
##  Prediction Diagnostics
cstat<- NULL
nll<- NULL
cstat<- apply(preds_all, 2, function(x) auc(dat\$A, x))   #auc
nll<- apply(preds_all, 2, function(x) nloglik(dat\$A, x)) #negative log-likelihood
##    Undersmoothing using balance diagnostics
bal_results1<- NULL
bal_results2<- NULL
bal_results3<- NULL
## undersmoothing based on iptw balance
bal_results1<- ps_undersmooth_bal(data=x_basis_full,
                                  treatment=dat\$A,
                                  ps_dat=preds_all,
                                  method=’iptw’,
                                  normalized=TRUE,
                                  standardize=TRUE)
## undersmoothing based on matching weight balance
bal_results2<- ps_undersmooth_bal(data=x_basis_full,
                                  treatment=dat\$A,
                                  ps_dat=preds_all,
                                  method=’mw’,
                                  normalized=TRUE,
                                  standardize=TRUE)
## undersmoothing based on overlap weight balance
bal_results3<- ps_undersmooth_bal(data=x_basis_full,
                                  treatment=dat\$A,
                                  ps_dat=preds_all,
                                  method=’ow’,
                                  normalized=TRUE,
                                  standardize=TRUE)
## cell with minimum ASAMD
bal_avg_counts1<- bal_results1\$index[1]
bal_avg_counts2<- bal_results2\$index[1]
bal_avg_counts3<- bal_results3\$index[1]
## cell with smallest maximum standardized difference
bal_max_counts1<- bal_results1\$index[2]
bal_max_counts2<- bal_results2\$index[2]
bal_max_counts3<- bal_results3\$index[2]
## asamd value
bal_avg1<- bal_results1\$balance[1]
bal_avg2<- bal_results2\$balance[1]
bal_avg3<- bal_results3\$balance[1]
## value of maximum standardized difference
bal_max1<- bal_results1\$balance[2]
bal_max2<- bal_results2\$balance[2]
bal_max3<- bal_results3\$balance[2]
bal_counts_avg<- c(bal_avg_counts1,
                   bal_avg_counts2,
                   bal_avg_counts3)
bal_counts_max<- c(bal_max_counts1,
                   bal_max_counts2,
                   bal_max_counts3)
bal_avg<- c(bal_avg1,
            bal_avg2,
            bal_avg3)
bal_max<- c(bal_max1,
            bal_max2,
            bal_max3)
names(bal_counts_avg)<- paste0(’count’, 1:length(bal_counts_avg))
names(bal_counts_max)<- paste0(’count’, 1:length(bal_counts_max))
names(bal_avg)<- paste0(’avg’, 1:length(bal_avg))
names(bal_max)<- paste0(’max’, 1:length(bal_max))
##  Estimating Treatment Effects
##   unadjusted estimate
est_crude <- NULL
est_crude <- mean(dat\$Y[dat\$A==1]) - mean(dat\$Y[dat\$A==0])
## adjusted estimates
est_ipw<- est_mw<- est_ow<- NULL
for(l in 1:ncol(preds_all)){
     A_hat<- preds_all[,l]
     ## IPW Estimation using Hajek estimator (normalized average, same as MLE)
     weight<- NULL
     weight<- dat\$A * 1/A_hat + (1-dat\$A) * 1/(1-A_hat)
     y1_ipw_est <- sum(dat\$A * dat\$Y * weight) / sum(dat\$A * weight)
     y0_ipw_est <- sum((1-dat\$A) * dat\$Y * weight) / sum((1-dat\$A) * weight)
     est_ipw[l] <- y1_ipw_est - y0_ipw_est
     ## Matching Weights using Hajek estimator (normalized average, same as MLE)
     score_data<- as.data.frame(cbind(A_hat, (1-A_hat)))
     numerator<- apply(score_data, 1, min)
     weight<- NULL
     weight<- numerator  / (dat\$A*A_hat + (1-dat\$A)*(1-A_hat))
     y1_mw_est<- sum(dat\$A * dat\$Y * weight) / sum(dat\$A * weight)
     y0_mw_est<- sum((1-dat\$A) * dat\$Y * weight) / sum((1-dat\$A) * weight)
     est_mw[l]<- y1_mw_est - y0_mw_est
     ## Overlap Weights using Hajek estimator (normalized average, same as MLE)
     weight<- NULL
     weight<- dat\$A * (1-A_hat) + (1-dat\$A) * A_hat
     y1_ow_est<- sum(dat\$A * dat\$Y * weight) / sum(dat\$A * weight)
     y0_ow_est<- sum((1-dat\$A) * dat\$Y * weight) / sum((1-dat\$A) * weight)
     est_ow[l]<- y1_ow_est - y0_ow_est
## Writing estimates to data frame
est_ipw_all<- as.data.frame(rbind(est_ipw))
est_mw_all<-  as.data.frame(rbind(est_mw))
est_ow_all<-  as.data.frame(rbind(est_ow))
colnames(est_ipw_all)<- paste0(’ipw_est’, 1:ncol(est_ipw_all))
colnames(est_mw_all)<-  paste0(’mw_est’,  1:ncol(est_mw_all))
colnames(est_ow_all)<-  paste0(’ow_est’,  1:ncol(est_ow_all))
## effect estimate for ipw, mw, and ow from cross-validated LASSO model
est_ipw_all[1]
est_mw_all[1]
est_ow_all[1]
## effect estimate for ipw, mw, and ow from undersmoothed LASSO model minimizing ASAMD
est_ipw_all[bal_avg_counts1]
est_mw_all[bal_avg_counts2]
est_ow_all[bal_avg_counts3]
## effect estimate for ipw, mw and ow from undersmoothed LASSO model with smallest maximum standardized difference
est_ipw_all[bal_max_counts1]
est_mw_all[bal_max_counts2]
est_ow_all[bal_max_counts3]
##  Generating Synthetic Negative Control Exposures and Running Analysis on Synthetic Cohorts
psS<- preds_all[,1] ##predicted values from CV Lasso fitted to full data (model that minimizes CV prediction error)
##  creating sampling probabilities
Xcovs_unexp <- x_basis_full[dat\$A==0,]                   # subsetting cohort to unexposed
prop_unexp <- psS[dat\$A==0]                              # subsetting predicted PS data to unexposed
theta<- log(prop_unexp/(1-prop_unexp))                   # log odds of the propensity score
exposureRate<- mean(dat\$A)                               # prevalence of exposure in the full population
fn <- function(c) mean(plogis(c + theta)) - exposureRate # function to find intercept (i.e., finds value for c so that function is 0)
delta <- uniroot(fn, lower = -100, upper = 100)\$root     # delta is intercept value
pi<- plogis(delta + theta)                               # pi are selection probabilities
y0<- dat\$Y[dat\$A==0]                                     # outcome in unexposed
##  Loop to generate and run analyses on many synthetic negative control exposure cohorts
nruns<- 500                                             #number of synthetic datasets to generate and run analyses on
est_ipw_synth_cv<- est_mw_synth_cv<- est_ow_synth_cv<- NULL #objects to store estimates from CV model
est_ipw_synth_b1<- est_mw_synth_b1<- est_ow_synth_b1<- NULL #objects to store estimates from undersmoothed model 1
est_ipw_synth_b2<- est_mw_synth_b2<- est_ow_synth_b2<- NULL #objects to store estimates from undersmoothed model 2
for(j in 1:nruns){
     ##  bootstrap oversampling and assigning synthetic exposure
     sample_index<- sample(1:nrow(Xcovs_unexp), nrow(x_basis_full), replace=TRUE) # sampling with replacement
     Xcovs_boot<- Xcovs_unexp[sample_index,]                                      # sample_index is the index for sampled individuals
     y0_boot<- y0[sample_index]                                                   # outcomes corresponding to sampled individuals
     pi_boot<- pi[sample_index]                                                   # probabilities for synthetic exposure
     zz<- rbinom(dim(Xcovs_boot)[1], 1, pi_boot)                                  # assignment of synthetic exposure status
     ## cleaning data to remove sparse variables
     temp1<- NULL
     temp2<- NULL
     temp1<- apply(Xcovs_boot, 2, mean)
     temp2<- (temp1 >= .01 & temp1 <= .99)
     Xcovs_boot<- Xcovs_boot[,temp2]
     ## refitting LASSO PS models in pseudo-population
     cvfolds<- NULL
     folds<- NULL
     foldid<- NULL
     cvfolds <- stratifyCVFoldsByYandID(V=nfolds, Y = zz)
     folds <- cvfolds\$validRows
     foldid <- cvfolds\$fold_id
     lasso_object<- treatment_model(data=Xcovs_boot,
                                    treatment=zz,
                                    foldid=foldid,
                                    alpha=1,
                                    lambda_ratio=0.01, #ifelse(nrow(x_basis_full) < ncol(x_basis_full), 0.01, 1e-04)
                                    nlambda=200,
                                    nmodels=NULL,
                                    maxit=5000,
                                    penalty=NULL,
                                    par=FALSE)
     preds<- NULL
     preds_cv<- NULL
     preds_all<- NULL
     coef_mat<- NULL
     lambdas<- NULL
     nvars<- NULL
     preds<- lasso_object[[1]]       #same-sample predictions
     preds_cv<- lasso_object[[2]]    #cross-validated (out-of-fold) predictions
     coef_mat<- lasso_object[[3]]    #coefficient matrix (coefficients for each lambda)
     lambdas<- lasso_object[[4]]     #lambda values
     nvars<- lasso_object[[5]]       #number of variables selected by each lambda
     #index values for subset of predictions and lambdas (select a range to reduce computation time instead of selecting all values)
     steps<- NULL
     steps<- floor(quantile(1:ncol(preds), seq(0, 1, .02)))
     preds_all<- preds_cv[,steps]
     coef_mat<- coef_mat[,steps]
     lambdas<- lambdas[steps]
     nvars<- nvars[steps]
     ## prediction diagnostics
     cstat<- NULL
     nll<- NULL
     cstat<- apply(preds_all, 2, function(x) auc(zz, x))
     nll<- apply(preds_all, 2, function(x) nloglik(zz, x))
     ## Undersmoothing using balance diagnostics
     bal_results1<- NULL
     bal_results2<- NULL
     bal_results3<- NULL
     bal_results1<- ps_undersmooth_bal(data=Xcovs_boot,
                                       treatment=zz,
                                       ps_dat=preds_all,
                                       method=’iptw’,
                                       normalized=TRUE,
                                       standardize=TRUE)
     bal_results2<- ps_undersmooth_bal(data=Xcovs_boot,
                                       treatment=zz,
                                       ps_dat=preds_all,
                                       method=’mw’,
                                       normalized=TRUE,
                                       standardize=TRUE)
     bal_results3<- ps_undersmooth_bal(data=Xcovs_boot,
                                       treatment=zz,
                                       ps_dat=preds_all,
                                       method=’ow’,
                                       normalized=TRUE,
                                       standardize=TRUE)
     ## cell with minimum ASAMD
     bal_avg_counts1<- bal_results1\$index[1]
     bal_avg_counts2<- bal_results2\$index[1]
     bal_avg_counts3<- bal_results3\$index[1]
     ## cell with smallest maximum standardized difference
     bal_max_counts1<- bal_results1\$index[2]
     bal_max_counts2<- bal_results2\$index[2]
     bal_max_counts3<- bal_results3\$index[2]
     ## asamd
     bal_avg1<- bal_results1\$balance[1]
     bal_avg2<- bal_results2\$balance[1]
     bal_avg3<- bal_results3\$balance[1]
     ## maximum standardized difference
     bal_max1<- bal_results1\$balance[2]
     bal_max2<- bal_results2\$balance[2]
     bal_max3<- bal_results3\$balance[2]
     bal_counts_avg<- c(bal_avg_counts1,
                        bal_avg_counts2,
                        bal_avg_counts3)
     bal_counts_max<- c(bal_max_counts1,
                        bal_max_counts2,
                        bal_max_counts3)
     bal_avg<- c(bal_avg1,
                 bal_avg2,
                 bal_avg3)
     bal_max<- c(bal_max1,
                 bal_max2,
                 bal_max3)
     names(bal_counts_avg)<- paste0(’count’, 1:length(bal_counts_avg))
     names(bal_counts_max)<- paste0(’count’, 1:length(bal_counts_max))
     names(bal_avg)<- paste0(’avg’, 1:length(bal_avg))
     names(bal_max)<- paste0(’max’, 1:length(bal_max))
     ##  Estimating Synthetic Exposure Effects
     ##   unadjusted synthetic estimate
     est_crude_synth<- NULL
     est_crude_synth<- mean(y0_boot[zz==1]) - mean(y0_boot[zz==0])
     ## adjusted synthetic estimates
     est_ipw_synth<- est_mw_synth<- est_ow_synth<- NULL
     for(l in 1:ncol(preds_all)){
          zz_hat<- preds_all[,l]
          ## IPW Estimation using Hajek estimator (normalized average, same as MLE)
          weight<- NULL
          weight<- zz * 1/zz_hat + (1-zz) * 1/(1-zz_hat)
          y1_ipw_synth <- sum(zz * y0_boot* weight) / sum(zz * weight)
          y0_ipw_synth <- sum((1-zz) * y0_boot * weight) / sum((1-zz) * weight)
          est_ipw_synth[l] <- y1_ipw_synth - y0_ipw_synth
          ## Matching Weights using Hajek estimator (normalized average, same as MLE)
          score_data<- as.data.frame(cbind(zz_hat, (1-zz_hat)))
          numerator<- apply(score_data, 1, min)
          weight<- NULL
          weight<- numerator  / (zz*zz_hat + (1-zz)*(1-zz_hat))
          y1_mw_synth<- sum(zz * y0_boot * weight) / sum(zz * weight)
          y0_mw_synth<- sum((1-zz) * y0_boot * weight) / sum((1-zz) * weight)
          est_mw_synth[l]<- y1_mw_synth - y0_mw_synth
          ## Overlap Weights using Hajek estimator (normalized average, same as MLE)
          weight<- NULL
          weight<- zz * (1-zz_hat) + (1-zz) * zz_hat
          y1_ow_synth<- sum(zz * y0_boot * weight) / sum(zz * weight)
          y0_ow_synth<- sum((1-zz) * y0_boot * weight) / sum((1-zz) * weight)
          est_ow_synth[l]<- y1_ow_synth - y0_ow_synth
     ## Writing estimates to data frame
     est_ipw_synth_all<- as.data.frame(rbind(est_ipw_synth))
     est_mw_synth_all<-  as.data.frame(rbind(est_mw_synth))
     est_ow_synth_all<-  as.data.frame(rbind(est_ow_synth))
     colnames(est_ipw_synth_all)<- paste0(’ipw_synth’, 1:ncol(est_ipw_synth_all))
     colnames(est_mw_synth_all)<-  paste0(’mw_synth’,  1:ncol(est_mw_synth_all))
     colnames(est_ow_synth_all)<-  paste0(’ow_synth’,  1:ncol(est_ow_synth_all))
     ## appending results for synthetic effect estimate for ipw, mw, and ow from cross-validated LASSO model
     est_ipw_synth_cv <- c(est_ipw_synth_cv, est_ipw_synth_all[1])
     est_mw_synth_cv <- c(est_mw_synth_cv, est_mw_synth_all[1])
     est_ow_synth_cv <-  c(est_ow_synth_cv, est_ow_synth_all[1])
     ## appending results for synthetic effect estimate for ipw, mw and ow from undersmoothed LASSO model minimizing ASAMD
     est_ipw_synth_b1 <- c(est_ipw_synth_b1, est_ipw_synth_all[bal_avg_counts1])
     est_mw_synth_b1 <- c(est_mw_synth_b1, est_mw_synth_all[bal_avg_counts2])
     est_ow_synth_b1 <- c(est_ow_synth_b1, est_ow_synth_all[bal_avg_counts3])
     ## appending results for synthetic effect estimate from undersmoothed LASSO model with smallest maximum standardized difference
     est_ipw_synth_b2 <- c(est_ipw_synth_b2, est_ipw_synth_all[bal_max_counts1])
     est_mw_synth_b2 <- c(est_mw_synth_b2, est_mw_synth_all[bal_max_counts2])
     est_ow_synth_b2 <- c(est_ow_synth_b2, est_ow_synth_all[bal_max_counts3])
## synthetic effect estimate for CV estimators averaged across all synthetic datasets
mean(as.numeric(est_ipw_synth_cv))
mean(as.numeric(est_mw_synth_cv))
mean(as.numeric(est_ow_synth_cv))
## synthetic effect estimate for undersmoothed estimators minimizing ASAMD averaged across all synthetic datasets
mean(as.numeric(est_ipw_synth_b1))
mean(as.numeric(est_mw_synth_b1))
mean(as.numeric(est_ow_synth_b1))
## synthetic effect estimate for undersmoothed estimators with smallest maximum st diff averaged across all synthetic datasets
mean(as.numeric(est_ipw_synth_b2))
mean(as.numeric(est_mw_synth_b2))
mean(as.numeric(est_ow_synth_b2))